# RuleRAG: Rule-Guided Retrieval-Augmented Generation with Language Models for Question Answering

### v1

@@ 5.2 Results of More LLMs
Figure 3: The Reacll@k and EM performance of RuleRAG-FT in RuleQA-I with different numbers of retrieved documents and under multiple circumstances: three settings in DPR (DPR, SSFT-DPR and RGFT-DPR), three settings in SimCSE (SimCSE, SSFT-SimCSE and RGFT-SimCSE) and one setting in BM25. The generator is kept as RGFT-LLAMA2_7B. Horizontal numbers over the pillars represent EM for bar charts and slanted numbers around the lines represent Recall@k for line charts.
5.3 Results of More Retrievers
5.3.1 More Retrievers for RuleRAG-FT
In Figure 3, we initialize RuleRAG-FT with more retrievers: dense retrievers DPR (Karpukhin et al., 2020), SimCSE (Gao et al., 2021) and training-free sparse retriever BM25 (Robertson & Zaragoza, 2009), and we use several retrieval configurations: retrievers without fine-tuning or with SSFT/RGFT while recalling different numbers of top-scored documents. Before fine-tuning, the Recall@k and EM performance of the three retrievers are comparable and each has its own performance, with no obvious advantages or disadvantages. For instance, comparing the three retrievers side-by-side, DPR has the best Recall@10 and SimCSE has the best EM under top-10 documents before fine-tuning.
After fine-tuning, DPR consistently outperforms SimCSE and RGFT consistently outperforms SSFT. Specifically, under considering top-scored documents with the same k, for the two trainable dense retrievers, the RGFT version recalls more relevant information (Recall@k) than the SSFT version by a large margin, demonstrating the generality of the proposed RGFT across different retrievers. As a result, the EM scores of the generated answers are better when higher-quality documents from retrievers are provided. Moreover, when the retrievers and generators are applied with RGFT, RuleRAG-FT shows substantial performance gains, even with the retrieval number limited to top-1. For DPR and SimCSE, as we include more documents, the Recall@k and EM scores increasingly improve. This shows that leveraging rules to guide the retrieval and generation processes builds a bridge between queries and answers since rules provide retrieval directions and attributable mechanisms. For BM25, although Recall@k keeps increasing, EM experiences a drop, probably due to the introduced noise.
One additional finding is that even though the difference in Recall@2 between the original DPR and SimCSE is not large (10.1% vs 10.2%), the EM of generated answers can differ significantly (12.4% vs 9.1%). The reason may be that the retrieved content of DPR includes not only the correct answers but also other helpful information. RGFT further widens the gap of Reacll@k between DPR and SimCSE.
5.3.2 More Retrievers for RuleRAG-ICL
Contriever (Izacard et al., 2022) is a powerful retriever with strong unsupervised performance and can transfer well to new applications. Therefore, it has been widely used in RAG frameworks. In Table 4, we note that Contriever without the guidance of rules can achieve relatively good recall and RG-Contriever makes further enhancements. Compared to Standard RAG, RuleRAG-ICL with RG-Contriever and RG-generators also obtain varying degrees of performance improvement under the three LLMs. These results confirm the outstanding ability of our proposed rule-guided method.
Table 4: The performance of RuleRAG-ICL with a powerful retriever, Contriever, under three LLMs.
Architecture RuleQA-I RuleQA-Y RuleQA-W RuleQA-F RuleQA-N
Figure 4: Due to the varying difficulty of the rules in five benchmarks, the EM performance variation of RuleRAG-FT produces different characteristics. The x-axis is the ratio of the amount of fine-tuned data to the total amount of fine-tuning data. The y-axis is the ratio of EM performance to the optimal one under DPR and LLAMA2_7B, with closer to 100% indicating stronger performance.
Figure 5: The EM performance of generalizing RuleRAG-FT from the source rule bank $\mathcal{R}_{i}$ to the target rule bank $\mathcal{R}_{j}$ , i.e., RuleRAG-FT is trained on $\mathcal{R}_{i}$ and tested on $\mathcal{R}_{j}$. The numbers in ($\mathcal{R}_{i}$, $\mathcal{R}_{j}$) represent the performance gains compared to the baseline Standard RAG tested on $\mathcal{R}_{j}$.
6 Impact of Fine-tuning Data Volume on EM Performance
In Figure 5, the overall performance trend of the five benchmarks is that the larger the amount of fine-tuning data, the better the results. Yet, since the different properties of the rules in different benchmarks lead to different degrees of difficulty in learning, the growth of model performance under different benchmarks exhibits various characteristics. Specifically, RuleQA-F has the most intuitive growth curve: It starts from low performance and slowly grows to the optimal performance, reflecting the increasing mastery of rules during the RGFT process of RuleRAG-FT. RuleQA-W yields relatively superior performance after just one-eighth of the total amount of the fine-tuning data and further improves subsequently. RuleQA-N achieves the best EM performance after three-eighths of the fine-tuning data and maintains flat. In contrast, the performance in RuleQA-Y fluctuates modestly at a very low level throughout the first half of the RGFT process (from one-eighth to five-eighths), and then sees a sudden surge in capability during the second half of the RGFT process (from six-eighths to the end). The EM performance in RuleQA-I fluctuates more dramatically: While realizing very large EM performance gains (ranking second in all the benchmarks), it undergoes several upward and downward drops before levelling off at the optimal performance. This suggests that RuleQA-I is the most challenging among our constructed five benchmarks. Moreover, from Table 2,  3,  4, we find RuleRAG has the worst absolute performance in RuleQA-I compared to other benchmarks under the same LLMs, which also illustrates the challenge of our constructed RuleQA-I.
7 Rule Generalization
RuleRAG-ICL is training-free, so we can attach arbitrary rules to the method’s input by in-context learning. Extensive experimental results above naturally illustrate its instruction-following ability to many kinds of rules. In the RGFT setting, the constructed fine-tuning data $\mathcal{F}_{G}$ for RuleRAG-FT is limited anyway but rules are inexhaustible, so RuleRAG-FT cannot and should not see the full set of rules. Therefore, it is important to verify the ability of RuleRAG-FT to generalize to untrained rules. In this experiment, RuleRAG-FT must capture the transferable rule utilization capability, since RuleRAG-FT has no prior knowledge of the target rule bank and is forced to learn from the source rule bank. The results in Figure 5, where $\mathcal{R}_{i}\cap\mathcal{R}_{j}=\emptyset$ and $\left|\mathcal{R}_{i}\right|=\left|\mathcal{R}_{j}\right|(i,j\in\{1,2,3,4\})$, show that (1) The diagonal ($\mathcal{R}_{i},\mathcal{R}_{i}$) has the highest performance gains and there are slight differences between various rule banks; (2) The results on two sides of the diagonal fluctuations within reasonable ranges and all show stable improvements over Standard RAG. This implies that RuleRAG-FT can take advantage of the ability to leverage the learned underlying rule patterns rather than being limited to concrete rule instances.
8 Related Works
8.1 Retrieval-Augmented Generation
Retrieval-augmented generation (RAG) follows the paradigm of “retrieve-then-read”, where the retrieval module explicitly augments the generation module with external knowledge banks (Guu et al., 2020; Lewis et al., 2020). Retrieval approaches include sparse retrievers based on sparse bag-of-words representation (Robertson & Zaragoza, 2009), dense retrievers based on dense vectors (Karpukhin et al., 2020; Gao et al., 2021) and more complex hybrid search algorithms (Li et al., 2023a; Lin et al., 2023). The current RAG frameworks are widely adopted to complement the parametric knowledge of LLMs along different stages (Gao et al., 2024), including pre-training stage (RETRO;  Borgeaud et al. (2022b), Atlas;  Izacard et al. (2024), COG;  Lan et al. (2023)), fine-tuning stage (Self-RAG;  Asai et al. (2023), SURGE;  Kang et al. (2023), CoN;  Yu et al. (2023)) and inference stage (DSP;  Khattab et al. (2023), KnowledGPT;  Wang et al. (2023), RoG;  Luo et al. (2024), CoK;  Li et al. (2024)).
8.2 Knowledge-intensive QA
In the realm of QA, a series of queries are considered knowledge-intensive if humans or models need access to large and external corpora. Researchers have developed many systems and proven the effectiveness of RAG in many knowledge-intensive QA tasks (Petroni et al., 2021). Recently, upon the assumption that documents in the corpora can directly support the answer responses, RAFT (Zhang et al., 2024) and RA-DIT (Lin et al., 2024) fine-tune LLMs by concatenating documents and queries as prompts. However, many answers to factual queries are hidden in semantically dissimilar but logically related events, which leads to irrelevant information retrieved by imperfect retrievers inevitably confusing RAG methods. Therefore, the integration of rules has gained significant attention (Wang et al., 2024b; c). For instance,  Wu et al. (2024) investigates the necessity of mitigating misleading irrelevant interference in RAG.  Sun et al. (2024) only discusses the rule-following abilities of LLMs without retrieval and ignores how to obtain these significant rules. In contrast, our proposed RuleRAG involves a more comprehensive consideration of mining rules, retrieving documents and generating answers.
9 Conclusion and Future Works
In this paper, we point out two high-level problems of current RAG and propose a method named rule-guided retrieval-augmented generation (RuleRAG) based on observations of the objective world. RuleRAG, including RuleRAG-ICL and RuleRAG-FT, can effectively improve the performance of multiple pre-trained retrievers and generators by in-context learning and instruction fine-tuning, respectively. RuleRAG-ICL intuitively shows that RAG can directly benefit from our proposal by prompting LLMs with rules. To further improve the QA performance, RuleRAG-FT retrofits the retrievers to recall more supportive information through the designed RGFT and updates generators to make better use of the retrieved documents. Experiments show RuleRAG achieves strong performance on the five constructed rule-aware QA benchmarks. In the future, we will explore how RuleRAG can effectively retrieve and answer when facing more complex queries and adapt to a wider range of rules.
Appendix A Appendix
A.1 The Robustness of RuleRAG
@@
Specifically, for each query in the benchmark, we degenerate it into a new relevant query by using the previously matched rules ( [Entity 1, $r_{1}$, Entity 2] leads to [Entity 1, $r_{2}$, Entity 2]) and ensure that the answer is unchanged and that the relevant documents can be retrieved directly from the corpus. Meanwhile, according to the principle of performance comparison, we try to minimize interference with the original queries. For instance in Figure 1, the original query is $\mathrm{What\ is\ the\ nationality\ of\ Jean\mbox{-}Luc\ Godard?}$ and the rule is that “ [Entity 1, born in, Entity 2] leads to [Entity 1, has nationality, Entity 2]”. Then, we convert the query into $\mathrm{Where\ is\ Jean\mbox{-}Luc\ Godard\ born?}$. In this way, these queries can theoretically be successfully retrieved with related documents and correctly answered without the guidance of rules.
@@
A.2 The Choice of RuleRAG-ICL and RuleRAG-FT
@@
A.3 The EM performance Trend of LLAMA2_7B and LLAMA2_13B
@@
A.4 RuleRAG-ICL with Contriever
As a complement to the performance of RuleRAG-ICL with Contriever in Table 4, we use the other two LLMs, Mistral_7B_v0.2 and LLAMA2_13B, to show the effectiveness of our proposed rule-guided method. Table 6 shows the performance of Mistral_7B_v0.2 and LLAMA2_13B with/without rules.
@@ A.4 RuleRAG-ICL with Contriever
A.5 Implementation Details
Generator fine-tuning. We fine-tune the ChatGLM2_6B, Mistral_7B_v0.2, LLAMA2_7B, LLAMA2_13B models using 2, 2, 4 and 8 V100 32G GPUs, respectively. We use LORA (Hu et al., 2022) with 4-bit, a parameter-efficient fine-tuning (PEFT) adaptation method, to deal with the enormous computation costs and hardware requirements in training LLM. The fine-tuning hyperparameters are detailed in Table 7. Similar to Lin et al. (2024), we find that the best generalization performance on the dev set can be achieved using a small number of fine-tuning epochs. We evaluate the models every 3 epochs and select the best checkpoint based on the average dev set performance.
Table 6: The results of RuleRAG-ICL with Contriever under Mistral_7B_v0.2 and LLAMA2_13B.
Architecture RuleQA-I RuleQA-Y RuleQA-W RuleQA-F RuleQA-N
RuleRAG-ICL RG-Contriever RG-Mistral_7B_v0.2 45.5 15.4 24.5 55.2 40.8 39.8 63.2 46.3 45.8 83.9 26.1 39.8 88.5 39.8 31.9
Standard RAG Contriever LLAMA2_13B 41.2 22.1 39.5 52.7 40.8 44.2 62.2 49.2 54.2 80.6 42.4 51.4 87.6 50.2 57.4
RuleRAG-ICL RG-Contriever RG-LLAMA2_13B 45.5 22.3 39.8 55.2 41.5 45.8 63.2 51.2 52.4 83.9 46.6 52.2 88.5 52.7 58.1
Table 7: Hyperparameters for RGFT-Generators.
ChatGLM2_6B 3e-5 4 16 0.05 5 8 50 1 5120
Mistral_7B_v0.2 3e-5 4 16 0.05 5 8 50 1 5120
LLAMA2_7B 3e-4 8 32 0.05 5 8 50 2 5120
LLAMA2_13B 3e-4 16 32 0.05 10 4 50 4 5120
Table 8: Instruct prompt.
# Query: Time 2014-12-01 what does Abdullah Abdullah Make a visit ?
# Answer: Afghanistan.
@@ A.5 Implementation Details
A.6 The Difference of Our Constructed Benchmarks and Existing Datasets
Existing conventional QA datasets include HotPotQA (Yang et al., 2018), 2WikiMultiopQA (Ho et al., 2020), MuSiQue (Trivedi et al., 2022), and Bamboogle (Press et al., 2023). Although they are widely leveraged in evaluating the QA performance of LMs, we find that all these datasets are primarily focused on multi-hop and comparison-type questions and pay less attention to queries that require logical thinking to answer. As we know, many queries in the real world are not justified by relevance alone, because in many cases the lexical level of relevance is not the information that can support the answer to the query, and even introduces a lot of noise instead. Therefore, in this paper, we construct five rule-aware benchmarks based on five popular static KGs or temporal KGs to emphasize the importance of rules in the QA task. It is worth noting that our described construction way in Section 2 is general and easy to reproduce. For newly defined rule patterns, we can quickly construct corresponding benchmarks using the above construction way, showing its better scalability.
Moreover, our constructed benchmarks also provide corresponding fine-tuning datasets, which aim to improve the retrieval and generation ability of LMs. Currently, obtaining high-quality and plentiful supervised data for a specific task is a challenging problem for researchers (Wang et al., 2024a). Manual annotation is time-consuming and difficult to replicate. A very convenient and widely used way is to distil knowledge from LLMs. However, relying on LLMs to generate data for training puts too much trust in them and does not actually guarantee the accuracy of the reasoning ability in the trained models.
@@ A.6 The Difference of Our Constructed Benchmarks and Existing Datasets
A.7 Case Study
@@ A.9 Prompt Templates
[…]

### v2

@@ Appendix C Further Analysis on LLMs
Appendix D Implementation Details
Generator fine-tuning. We fine-tune the ChatGLM2_6B, Mistral_7B_v0.2, LLAMA2_7B, LLAMA2_13B models using 2, 2, 4 and 8 V100 32G GPUs, respectively. We use LORA (Hu et al., 2022) with 4-bit, a parameter-efficient fine-tuning (PEFT) adaptation method, to deal with the enormous computation costs and hardware requirements in training LLM. Hyper-parameter N is 3 and $\theta$ is 0.7. The fine-tuning hyperparameters are detailed in Table 6. Similar to Lin et al. (2024), we find that the best generalization performance on the dev set can be achieved using a small number of fine-tuning epochs. We evaluate the models every 3 epochs and select the best checkpoint based on the average dev set performance.
ChatGLM2_6B 3e-5 4 16 0.05 5 8 50 1 5120
Mistral_7B_v0.2 3e-5 4 16 0.05 5 8 50 1 5120
LLAMA2_7B 3e-4 8 32 0.05 5 8 50 2 5120
LLAMA2_13B 3e-4 16 32 0.05 10 4 50 4 5120
Table 6: Hyperparameters for RGFT-Generators.
Appendix E The Robustness of RuleRAG
@@
Specifically, for each query in the benchmark, we degenerate it into a new relevant query by using the previously matched rules ( [Entity 1, $r_{1}$, Entity 2] leads to [Entity 1, $r_{2}$, Entity 2]) and ensure that the answer is unchanged and that the relevant documents can be retrieved directly from the corpus. Meanwhile, according to the principle of performance comparison, we try to minimize interference with the original queries. For instance, the original query is $\mathrm{What\ is\ the\ nationality\ of\ Jean\mbox{-}Luc\ Godard?}$ and the rule is that “ [Entity 1, born in, Entity 2] leads to [Entity 1, has nationality, Entity 2]”. Then, we convert the query into $\mathrm{Where\ is\ Jean\mbox{-}Luc\ Godard\ born?}$. In this way, these queries can theoretically be successfully retrieved with related documents and correctly answered without the guidance of rules.
@@
To further improve the robustness of RuleRAG, in future work, we can use LLM to filter, sort, and evaluate rules or consider rules as interactable logical units, and so on. For exceptions or anomalies, we can also introduce entity linking for unrecognized entities and semantic similarity checks for outliers in temporal data. In addition, the robustness of the LLM itself can also ensure performance.
@@
Appendix F The Choice of RuleRAG-ICL and RuleRAG-FT
@@
Appendix G The EM performance Trend of LLAMA2_7B and LLAMA2_13B
@@
Appendix H The Difference of RuleQA and Existing RAG Datasets
Most existing RAG datasets for QA only provide questions and corpora when construction and do not match suitable rules for them. In this paper, we construct five rule-aware QA benchmarks RuleQA, where many high-quality rules are mined from KGs to guide the retrieval and reasoning. Meanwhile, we experimentally show that both the introduced rules and our proposed model RuleRAG are effective in four existing RAG datasets, ASQA (Stelmakh et al., 2022) (long-form QA), PopQA (Mallen et al., 2023)(short-form QA), HotpotQA (Yang et al., 2018)(multi-hop QA) and Natural Questions (NQ) (Kwiatkowski et al., 2019).
In addition, although they are widely leveraged in evaluating the QA performance of LMs, we find that all these datasets are primarily focused on multi-hop and comparison-type questions and pay less attention to queries that require logical thinking to reason. As we know, many queries in the real world are not justified by relevance alone, because in many cases the lexical level of relevance is not the information that can support the answer to the query, and even introduces a lot of noise instead. Therefore, in this paper, we construct five rule-aware QA benchmarks RuleQA based on five popular static KGs or temporal KGs to emphasize the importance of rules in the QA task. It is worth noting that our described construction way in Section A is general and easy to reproduce. For newly defined rule patterns, we can quickly construct corresponding benchmarks using the above construction way, showing its better scalability.
Moreover, our constructed RuleQA also provide corresponding fine-tuning datasets, which aim to improve the retrieval and generation ability of LMs. Currently, obtaining high-quality and plentiful supervised data for a specific task is a challenging problem for researchers (Wang et al., 2024). Manual annotation is time-consuming and difficult to replicate. A very convenient and widely used way is to distil knowledge from LLMs. However, relying on LLMs to generate data for training puts too much trust in them and does not actually guarantee the accuracy of the reasoning ability in the trained models.
@@ Appendix H The Difference of RuleQA and Existing RAG Datasets
# Query: Time 2014-12-01 what does Abdullah Abdullah Make a visit ?
# Answer: Afghanistan.
Table 9: Instruct prompt.
Appendix I Case Study
@@
# Retrieved documents: Documents related to the Query. Time 2014-03-11 Alexis Tsipras Make a visit Ireland. Time 2014-02-26 Alexis Tsipras Express intent to meet or negotiate Slovenia. $\cdots$ Time 2014-05-26 Alexis Tsipras Make a visit Head of Government (Greece). $\cdots$ Time 2014-09-17 Alexis Tsipras Consult New Democracy.